a certain quantity of agas occupies 200ml when collected over water at 15 degree temperature(c) and 745mm pressure.it occupies 182.6ml in dry state at NTP.find thevpourpressure of water at 15(degree centigrade)
 
 
 
another question
calculate he volume occupied by 5g of acetylene gas at 50 degree centigrade nd 740mm of pressure
 

Grade:12

2 Answers

Arun
4 years ago
Applying the general gas equation
PV = nRT = m/M RT
Here, Mol. wt. of acetylene i.e., C2H2 (M) = 26, P = 740/760 atm,
T = 50oC = 50 + 273 = 323 K
∴ V = mRT/MP or V 5 * 0.082 *323 *760/26 *740 = 5.23 l
Vikas TU
4 years ago
Dear student 
Please ask only single question in a thread, It will be easier fr us to ans.
Hope you willunderstand 
Good Luck 
Cheers